OpenStudy (anonymous):

Is my answer right? Problem: 2x^2+8x+8/2x^2+2x-4 My answer: (x+4)/(x-2)

OpenStudy (phi):

I think you went wrong somewheres...

OpenStudy (anonymous):

wants me to simplify

OpenStudy (phi):

assuming the quadratics are over each other

OpenStudy (hoblos):

2x^2+8x+8 = 2(x+2)^2 2x^2+2x-4 = 2(x+2)(x-1) now try to simplify

OpenStudy (phi):

You should first factor a 2 out of the top and a 2 out of the bottom and cancel them

OpenStudy (phi):

see hoblos post

OpenStudy (phi):

then factor the quadratics x^2+4x+4 and x^2+x-2

OpenStudy (phi):

Do you see where you went wrong?

OpenStudy (anonymous):

yaaaaaaaaa its wrong

OpenStudy (anonymous):

(x+2)/(x-1) this right? and yes i think so

OpenStudy (phi):

absolutely

OpenStudy (hoblos):

yeah now you are right :)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

:D thanks guys!